OpenStudy (youngnephew34):

The measure of two complementary angles are in the ratio 2:3. What is the measurement of the smaller angle

OpenStudy (youngnephew34):

I have a question bout this problem

OpenStudy (karolynxx):

Step 1. Let 2x be one angle and 3x be the other angle. The ratio of these two angles is 2x/3x=2/3 as given on the problem. Step 2. The two angles must add up to 180 degrees by definition. Step 3. Then 2x+3x=180 or 5x=180 which implies that x=36. 2x=72 and 3x=108 Step 4. The angles are then 72 degrees and 108 degrees. Note the two angles add up to 180 and their ratio is 72/108=2/3 satisfying the problem statement.

OpenStudy (youngnephew34):

yea but I only had a question

OpenStudy (youngnephew34):

don't you suppose to do 2x3 but nvm

OpenStudy (karolynxx):

Oh sorry , what is it?

OpenStudy (youngnephew34):

nvm nvm